OpenStudy (anonymous):

The probability of pulling a black cap out of a cupboard is 1/x.the probability of pulling a black cap out of a cupboard and dropping a glove is 1/y.what is the probability of dropping a glove? 1/(xy), x/y, y/x, 1/x+y

OpenStudy (anonymous):

help me

OpenStudy (jamesj):

In general, if A and B are events, Pr(A and B) = Pr(A|B) . Pr(B) For Pr(A|B) we say "The probability of A given B"

OpenStudy (anonymous):

thks for the help jemesj

OpenStudy (anonymous):

can u solve this also A rectangular field is divided in to a metre grid like in the coordinate system .centre being the origin of the grid. What is the area of the rectangular flower bed with the following coordinates (x,y),(10,y),(10,5),(x,5) in the field? 6, 8, 12, cannot be determined

OpenStudy (jamesj):

On your original question Pr(black cap) = 1/x Pr(black cap and dropped glove) = 1/y Now it's also true if the events are independent that Pr(black cap and dropped glove) = Pr(black cap).Pr(dropped glove) Hence Pr(dropped glove) = Pr(black cap and dropped glove) / Pr(black cap) = 1/y / (1/x) = x/y

OpenStudy (jamesj):

Post your new question on the left and someone will help you.